OpenStudy (anonymous):

What is the side length of a cube with a volume of 2197 cm3?

OpenStudy (anonymous):

Volume of a cube = (side)^3 So solve 2197 = (side)^3.

OpenStudy (anonymous):

side will be cube root of 2197.

OpenStudy (anonymous):

answer

OpenStudy (wolf1728):

Studyislandhelp Don't you know how to calculate a cube root?

OpenStudy (charlotte123):

trying to solve this >.> sorry for taking a lot of time, my internet isn't coping .-.

OpenStudy (anonymous):

13

OpenStudy (anonymous):

wats the answer